OpenStudy (anonymous):

Solving Exponential Equations with Unequal Bases Problem: 3^(x-4)=7^(x+9) Answer Options: x ≈ –25.86 x ≈ –0.04 x ≈ –0.39 x ≈ –2.59

OpenStudy (experimentx):

take log on both sides \( \ln3^{x-4} = \ln{7^{x+9}}\) \( = (x-4)\ln 3 = (x+9)\ln7\) <--- solve this equation.

OpenStudy (anonymous):

i dont know how to solve it, i have no idea how to work these problems.

OpenStudy (experimentx):

use calculator to calculate the value of log. http://www.wolframalpha.com/input/?i=solve+%28x%E2%88%924%29ln3%3D%28x%2B9%29ln7